What is the tangential acceleration of a bug on the rim of a 9.0 inch diameter disk if the disk moves from rest to an angular speed of 90 revolutions per minute in 4.0 seconds?


1
Expert's answer
2021-12-12T16:39:58-0500

The tangential acceleration is


"a=r\alpha=r\frac{2\pi n }{t}=21.2\text{ in/s}^2."